OpenStudy (anonymous):

What is the value of y in the equation 4(5y − 8 − 2) = 185 − 15? 8.5 9.7 10.5 42.5

OpenStudy (study_together):

20y-32-8=170 20y-40=170 +40 +40 210 20y/20 210/20 Answer=10.5 Answer=C